# Labeling eigenstates of qubit-cavity systems based on the continuity of qubit occupancy: Detecting resonances to higher excited qubit states

### v1

@@ I Introduction
The rest of the paper is organized as follows: In Sec. II, we introduce labeling methods used in this paper. The proposed approach is given in this section. In Sec. III, we compare the existing and the proposed labeling methods. In Sec. IV, the offset charge dependence of resonances to higher excited qubit states is investigated with the proposed method. The summary is given in Sec. V.
@@ II Labeling methods
We discuss how to find a labeled state for a qubit state |$\ket{p}_{q}$ and a Fock state |$\ket{n}_{c}$, |$\ket{\overline{p},\overline{n}}$, from the eigenstates of $\hat{H}$. Starting from an initial state |$\ket{\overline{p},\overline{0}}$, the system is expected to follow the labeled states |$\ket{\overline{p},\overline{n}}$ under adiabatic injection of photons into the cavity [26, 22, 21]. The eigenenergy of the labeled eigenstate |$\ket{\overline{p},\overline{n}}$ is denoted by $\varepsilon_{p,n}$. With the labeled eigenenergies, one can regard $(\varepsilon_{p,n+1}-\varepsilon_{p,n})/\hbar$ as the effective cavity frequency when the qubit is initially |$\ket{p}_{q}$ and the cavity photon number is $n$. The photon-number dependence of the cavity frequency is useful information for designing the parameters of the dispersive readout [21, 22]. Besides, a resonance to a higher excited qubit state becomes visible in the photon-number dependence of the qubit frequency.
@@
$\varepsilon^{\prime}_{p,n}=\varepsilon_{p,n-1}+\hbar\omega_{c}.$ (3)
Next, one selects every eigenstate whose eigenenergy $\varepsilon$ satisfies $|\varepsilon-\varepsilon^{\prime}_{p,n}|\leq\delta/2$, where $\delta$ is the energy window. The parameter $\delta$ should be large enough to detect sharp peaks coming from resonances and small enough not to include unnecessary states. We set the parameter $\delta$ to $1.0\times 10^{-2}\hbar\omega_{c}$ in this paper. If no eigenstate is found within the energy window, one selects the two eigenstates closest to the candidate. For the selected eigenstates, one evaluates the expectation values of $\hat{N}_{q}$. The next labeled state |$\ket{\overline{p},\overline{n}}$ is determined as the state with the closest expectation value to $\braket{\overline{p},\overline{n-1}}{\hat{N}_{q}}{\overline{p},\overline{n-1}}$. In short, the proposed method selects the state that makes the change of qubit occupancy as small as possible.
@@ III Comparison of labeling methods
Figure 1: Photon-number dependence of the cavity frequencies in the transmon-cavity system obtained by the three labeling methods. The recursive calculations start from |$\ket{\overline{g},\overline{0}}$. For the system parameters, we use $(E_{C}/\hbar\omega_{c},E_{J}/\hbar\omega_{c},g/\omega_{c},N_{g})=(5.0\times 10^{-2},1.6,2.5\times 10^{-2},0.0)$.
Figure 1 gives the photon-number dependence of the cavity frequencies obtained by the three labeling methods. The recursive calculations start from |$\ket{\overline{g},\overline{0}}$. We use $(E_{C}/\hbar\omega_{c},E_{J}/\hbar\omega_{c},g/\omega_{c},N_{g})=(5.0\times 10^{-2},1.6,2.5\times 10^{-2},0.0)$ for the system parameters. The three labeling methods result in the same labeling up to $n\sim 150$. The sharp peak around $n\sim 50$ comes from the resonance with the higher excited qubit states. Around $n\sim 150$, the frequency given by the overlap method gives a discontinuous jump. After the discontinuous jump, the cavity frequency obtained by the overlap method shows non-smooth dependence. The observation of unphysical non-smooth behavior means that the overlap method fails to trace the labeled states |$\ket{\overline{g},\overline{n}}$, and the failure implies that the labeled states |$\ket{\overline{g},\overline{n}}$ are not close to |$\ket{g}_{q}\otimes\ket{n}_{c}$ for $n\gtrsim 150$.
@@
where $\braket{\hat{A}(t)}_{U}$ stands for $\braket{\psi(t)}{\hat{U}^{\dagger}(t)\hat{A}\hat{U}(t)}{\psi(t)}$ [22]. For the numerical simulation of the dynamics, we use the Fock state of the cavity up to |$\ket{150}_{c}$, and the drive frequency is set to $\omega_{d}/\omega_{c}=1.0015$.
Figure 2: Photon-number dependence of the qubit occupation number in the transmon-cavity system obtained by the two labeling methods (recursive and proposed) and the evolution under cavity driving. The recursive calculations start from |$\ket{\overline{g},\overline{0}}$. An initial state for the dynamics is also set to |$\ket{\overline{g},\overline{0}}$. The parameters of the systems are the same with those in Fig. 1. The amplitude of the driving field $E$ is set to $5.0\times 10^{-3}\omega_{c}$, and the drive frequency $\omega_{d}$ is set to $1.0015\omega_{c}$.
Figure 2 gives the photon-number dependence of the qubit occupation number obtained by the recursive and the proposed methods. The evolution of the qubit occupation number under the cavity driving is also plotted in Fig. 2. The qubit occupation numbers obtained from the two labeling methods show different behaviors for $n\gtrsim 180$, and their behaviors are very similar to those in the cavity frequency shown in Fig. 1. The recursive method gives discontinuous increase of the qubit occupation number at $n\sim 180$. On the contrary, that obtained by the proposed method exhibits a sharp peak at the point.
Under cavity driving, the evolution of the qubit occupation number roughly follows the photon-number dependence obtained by the proposed method apart from the resonant peaks. Therefore, we conclude that the proposed method can be applied to a broader parameter region compared to the recursive method. We present another numerical example with parameters $(E_{C}/\hbar\omega_{c},E_{J}/\hbar\omega_{c},g/\omega_{c},N_{g},\delta/\hbar\omega_{c},E/\omega_{c},\omega_{d}/\omega_{c})=(5.0\times 10^{-2},2.0,1.0\times 10^{-2},0.1,1.0\times 10^{-2},5.0\times 10^{-3},1.0002)$, and the same conclusion is obtained as summarized in Fig. 3. The recursive method fails to follow the evolution of states at the resonant point, while the proposed method succeeds.
@@
IV Offset charge dependence of resonance
@@
Because of the infinite sum in the charge operator defined in Eq. (5), the operators before and after the shift of the offset charge by unity, $N_{g}\to N_{g}+1$, are unitary equivalent, i.e.,
@@
with the unitary transformation $\hat{V}_{\mathrm{invert}}=\sum\ket{n}_{t}\bra{-n}_{t}$. The unitary transformation $\hat{V}_{\mathrm{invert}}$ does not change the Josephson energy term, neither. Since the negative sign can be absorbed into the definition of the cavity annihilation and creation operators, the spectrum of the transmon-cavity system is invariant under the sign inversion of the offset charge. From the periodicity and the symmetry, it is enough to investigate the region $0\leq N_{g}\leq 0.5$ for the identification.
@@
Figure 4: Photon-number dependence of the cavity frequency with the offset charges from 00 to $0.5$ for (a) the ground-state ladder|$\ket{\overline{g},\overline{n}}$ and (b) the excited-state ladder |$\ket{\overline{e},\overline{n}}$. We use $(E_{C}/\hbar\omega_{c},E_{J}/\hbar\omega_{c},g/\omega_{c},N_{g})=(5.0\times 10^{-2},1.6,2.5\times 10^{-2})$ for the system parameters. The energy window used for the labeling is $\delta/\hbar\omega_{c}=1.0\times 10^{-2}$.
@@
Figure 4 shows the photon-number dependence of the cavity frequency with the offset charges from 0 to 0.5 for the ground- (excited-)state ladder |$\ket{\overline{g},\overline{n}}$ (|$\ket{\overline{e},\overline{n}}$). For the ground-state ladder |$\ket{\overline{g},\overline{n}}$, the cavity frequencies for the different offset charges do not show visible difference up to $n\sim 50$ apart from the resonant peaks. The little dependence on the offset charge implies that the weights of higher transmon excited states in the labeled states |$\ket{\overline{g},\overline{n}}$ are small for $n\lesssim 50$ (apart from the resonant point). The location of the resonant point strongly depends on the offset charge, as expected. The resonant point can fluctuate between $n\sim 15$ and $n\sim 50$ without controlling the offset charge. Therefore, a transition to a higher excited state may take place with a relatively small photon number $n\sim O(10)$.
@@
For the excited-state ladder |$\ket{\overline{e},\overline{n}}$, the cavity frequencies show visible difference for $n\gtrsim 20$. This implies that the weights of higher transmon excited states are not negligible in this region. The resonant peak is present only in the case with $N_{g}=0.5$, and the location of the peak is around $n\sim 10$. Consequently, a transition to a higher excited state may take place with a relatively small photon number also in the case of the labeled states |$\ket{\overline{e},\overline{n}}$. In this way, the photon-number dependence of the cavity frequency obtained by the labeling method provides the information on photon number where the transitions leading to leakage from the computational space can occur.
V Summary
@@
In this paper, we proposed a new method to label the eigenstates of a qubit-cavity system based on the continuity of qubit occupancy. Starting from the eigenstate without photon occupation, the system is expected to follow the labeled states under adiabatic injection of photons into the cavity. Besides, the effective resonant frequency of the cavity with photon occupation can be estimated from the eigenenergies of the labeled states. Our proposed method finds candidates for the next labeled state based on their eigenenergies. Then, one selects the next labeled state that makes the change of the qubit occupancy as small as possible. Comparing the obtained photon-number dependence of the cavity frequency and the evolution under cavity driving, we confirmed that our proposed scheme can be applied to broader systems compared to the previous approach [21].

### v2

@@ I Introduction
The rest of the paper is organized as follows: In Sec. II, we introduce labeling methods used in this paper. The proposed approach is given in this section. In Sec. III, we compare the existing and the proposed labeling methods. The behavior of the proposed method can be controlled by the energy window. We discuss how to choose an appropriate energy window in section IV. In Sec. V, the offset charge dependence of resonances to higher excited qubit states is investigated with the proposed method. The summary is given in Sec. VI.
@@ II Labeling methods
We discuss how to find a labeled state for a qubit state |$\ket{p}_{q}$ and a Fock state |$\ket{n}_{c}$, |$\ket{\overline{p},\overline{n}}$, from the eigenstates of $\hat{H}$. Starting from an initial state |$\ket{\overline{p},\overline{0}}$, the system is expected to follow the labeled states |$\ket{\overline{p},\overline{n}}$ under adiabatic injection of photons into the cavity [26, 22, 21]. The eigenenergy of the labeled eigenstate |$\ket{\overline{p},\overline{n}}$ is denoted by $\varepsilon_{p,n}$. With the labeled eigenenergies, one can regard $(\varepsilon_{p,n+1}-\varepsilon_{p,n})/\hbar$ as the effective cavity frequency when the qubit is initially |$\ket{p}_{q}$ and the cavity photon number is $n$. The photon-number dependence of the cavity frequency is useful information for designing the parameters of the dispersive readout [21, 22]. Besides, a resonance to a higher excited qubit state becomes visible in the photon-number dependence of the cavity frequency.
@@
$\varepsilon^{\prime}_{p,n}=\varepsilon_{p,n-1}+(\varepsilon_{p,n-1}-\varepsilon_{p,n-2})$ (3)
for $n\geq 2$ and
$\varepsilon^{\prime}_{p,1}=\varepsilon_{p,0}+\hbar\omega_{c}$ (4)
for $n=1$. The quantity $\varepsilon_{p,n-1}-\varepsilon_{p,n-2}$ corresponds to the effective energy shift induced by adding a single photon. Next, one selects every eigenstate whose eigenenergy $\varepsilon$ satisfies $|\varepsilon-\varepsilon^{\prime}_{p,n}|\leq\delta/2$, where $\delta$ is the energy window. The parameter $\delta$ should be large enough to detect sharp peaks coming from resonances and small enough not to include unnecessary states. If no eigenstate is found within the energy window, one selects the two eigenstates closest to the candidate. For the selected eigenstates, one evaluates the expectation values of $\hat{N}_{q}$. The next labeled state |$\ket{\overline{p},\overline{n}}$ is determined as the state with the closest expectation value to $\braket{\overline{p},\overline{n-1}}{\hat{N}_{q}}{\overline{p},\overline{n-1}}$. In determining |$\ket{\bar{p},\bar{1}}$, the overlap approach is also available. In short, the proposed method selects the state that makes the change of qubit occupancy as small as possible.
@@ III Comparison of labeling methods
Figure 1: Photon-number dependence of the cavity frequencies in the transmon-cavity system obtained by the three labeling methods. The recursive calculations start from |$\ket{\overline{g},\overline{0}}$. For the system parameters, we use $(E_{C}/\hbar\omega_{c},E_{J}/\hbar\omega_{c},g/\omega_{c},N_{g})=(5.0\times 10^{-2},1.6,2.5\times 10^{-2},0.0)$. The energy window $\delta$ is set to $1.0\times 10^{-2}\hbar\omega_{c}$ for the proposed method.
Figure 1 gives the photon-number dependence of the cavity frequencies obtained by the three labeling methods. The recursive calculations start from |$\ket{\overline{g},\overline{0}}$. We use $(E_{C}/\hbar\omega_{c},E_{J}/\hbar\omega_{c},g/\omega_{c},N_{g})=(5.0\times 10^{-2},1.6,2.5\times 10^{-2},0.0)$ for the system parameters. The energy window is set to $\delta=1.0\times 10^{-2}\hbar\omega_{c}$. The three labeling methods result in the same labeling up to $n\sim 150$. The sharp peak around $n\sim 50$ comes from the resonance with the higher excited qubit states. The transition to higher excited state occurs when the cavity photon number stays near the resonant peak. However, such transition can be avoided by quickly passing the resonant point [21, 22]. Around $n\sim 150$, the frequency given by the overlap method gives a discontinuous jump. After the discontinuous jump, the cavity frequency obtained by the overlap method shows non-smooth dependence. The observation of unphysical non-smooth behavior means that the overlap method fails to trace the labeled states |$\ket{\overline{g},\overline{n}}$, and the failure implies that the labeled states |$\ket{\overline{g},\overline{n}}$ are not close to |$\ket{g}_{q}\otimes\ket{n}_{c}$ for $n\gtrsim 150$.
@@
where $\braket{\hat{A}(t)}_{U}$ stands for $\braket{\psi(t)}{\hat{U}^{\dagger}(t)\hat{A}\hat{U}(t)}{\psi(t)}$ [22]. For the numerical simulation of the dynamics in this section, we use the Fock state of the cavity up to |$\ket{150}_{c}$, and the drive frequency is set to $\omega_{d}/\omega_{c}=1.0015$.
Figure 2: Photon-number dependence of the qubit occupation number in the transmon-cavity system obtained by the two labeling methods (recursive and proposed) and the evolution under cavity driving. The recursive calculations start from |$\ket{\overline{g},\overline{0}}$. An initial state for the dynamics is also set to |$\ket{\overline{g},\overline{0}}$. The parameters of the system are the same as those in Fig. 1. The amplitude of the driving field $E$ is set to $5.0\times 10^{-3}\omega_{c}$, and the drive frequency $\omega_{d}$ is set to $1.0015\omega_{c}$.
Figure 2 gives the photon-number dependence of the qubit occupation number obtained by the recursive and the proposed methods. We also plot the trajectory in the $\braket{\hat{N}_{q}}$-$\braket{\hat{c}^{\dagger}\hat{c}}$ plane under the cavity driving with $E/\omega_{c}=5.0\times 10^{-3}$in Fig. 2. The amplitude of the drive field $E$ is determined strong enough not to be affected by resonances from higher excited states. The qubit occupation numbers obtained from the two labeling methods show different behaviors for $n\gtrsim 180$, and their behaviors are very similar to those in the cavity frequency shown in Fig. 1. The recursive method gives discontinuous increase of the qubit occupation number at $n\sim 180$. On the contrary, that obtained by the proposed method exhibits a sharp peak at the point.
Under cavity driving, the evolution of the qubit occupation number roughly follows the photon-number dependence obtained by the proposed method apart from the resonant peaks. Therefore, we conclude that the proposed method can be applied to a broader parameter region compared to the recursive method. We present another numerical example with parameters $(E_{C}/\hbar\omega_{c},E_{J}/\hbar\omega_{c},g/\omega_{c},N_{g},\delta/\hbar\omega_{c},E/\omega_{c},\omega_{d}/\omega_{c})=(5.0\times 10^{-2},2.0,1.0\times 10^{-2},0.1,1.0\times 10^{-2},5.0\times 10^{-3},1.0002)$ starting from the excited state |$\ket{\bar{e},\bar{0}}$. For this case with a finite charge offset, the same conclusion is obtained as summarized in Fig. 3. The recursive method fails to follow the evolution of states at the resonant point, while the proposed method succeeds. The proposed method can access the effective cavity frequency beyond the resonant peak.
@@
IV How to choose appropriate energy window
Figure 4: Photon-number dependence of (a) the cavity frequency and (b) the qubit occupation number in the transmon-cavity system obtained by the proposed labeling methods with different energy windows $\delta$. For the qubit occupation number, the evolution under cavity driving is also plotted. The labeling method starts from |$\ket{\overline{g},\overline{0}}$. An initial state for the dynamics is also set to |$\ket{\overline{g},\overline{0}}$. For the calculation, we use $(E_{C}/\hbar\omega_{c},E_{J}/\hbar\omega_{c},g/\omega_{c},N_{g},E/\omega_{c},\omega_{d}/\omega_{c})=(5.0\times 10^{-2},1.6,2.5\times 10^{-2},0.1,1.0\times 10^{-2},1.001)$.
Compared to the recursive approach, our proposed method introduces an additional parameter, the energy window $\delta$. This parameter should be large enough to include the eigenenergy affected by a higher excited state within the energy window and small enough to exclude eigenstates whose qubit occupancies are accidentally close to the target value. The additional parameter can adjust the behavior of the labeling method, and thus the proposed method acquires adaptability. In this section, we discuss how to choose an appropriate energy window $\delta$.
To discuss the appropriate choice of the energy window, we observe the photon-number dependence of the effective cavity frequency and the qubit occupancy given by the proposed labeling method with different energy windows, as shown in Fig. 4. We use the system parameters $(E_{C}/\hbar\omega_{c},E_{J}/\hbar\omega_{c},g/\omega_{c},N_{g})=(5.0\times 10^{-2},1.6,2.5\times 10^{-2},0.1)$ and three distinct energy windows $1.0\times 10^{-2}\hbar\omega_{c}$, $1.5\times 10^{-2}\hbar\omega_{c}$, and $2.0\times 10^{-2}\hbar\omega_{c}$. For the largest energy window $2.0\times 10^{-2}\hbar\omega_{c}$, even the effective cavity frequency shows a discontinuous jump, and this jump suggests that the selected energy window is too large and includes an unnecessary state. For the other two energy windows, a discontinuous behavior is not visible in the cavity frequency and the qubit occupancy except for resonant peaks.
By recalling the results presented in Sec. III, a resonant peak of the cavity frequency is replaced with a sudden drop in failed labeling. Based on this empirical rule, the labeling with the energy window $\delta=1.5\times 10^{-2}\hbar\omega_{c}$ should be appropriate. This expectation can be verified by comparing the trajectory obtained by the dynamics under cavity driving. The red-dotted line in Fig. 4(b) is the trajectory in the $\braket{\hat{N}_{q}}$-$\braket{\hat{c}^{\dagger}\hat{c}}$ plane under the cavity driving with $E/\omega_{c}=1.0\times 10^{-2}$ and $\omega_{d}/\omega_{c}=1.001$. We use the Fock state of the cavity up to |$\ket{500}_{c}$ for the numerical simulation of the dynamics. The trajectory obtained by the dynamics follows the photon number dependence of the qubit occupancy with the energy window $\delta/\hbar\omega_{c}=1.5\times 10^{-2}$. The dynamics ensure that the labeling with $\delta/\hbar\omega_{c}=1.5\times 10^{-2}$ is appropriate. This rough matching with the dynamics also implies that a drop of the cavity frequency around $n\sim 240$ will not be replaced with a resonant peak.
Through this example, one can find the importance of using some energy windows and comparing those results. From the obtained results, one can eliminate labeling with an evident discontinuity in physical quantities. When a sudden change in the cavity frequency is replaced with a peak in another labeling, the labeling with the sudden change would be wrong. This empirical rule is useful for choosing an appropriate energy window without accessing dynamics. For the numerically expensive but definite confirmation, one can compare the results of labeling to those obtained from dynamics.
V Offset charge dependence of resonance
@@
Because of the infinite sum in the charge operator defined in Eq. (6), the operators before and after the shift of the offset charge by unity, $N_{g}\to N_{g}+1$, are unitarily equivalent, i.e.,
@@
with the unitary transformation $\hat{V}_{\mathrm{invert}}=\sum\ket{n}_{t}\bra{-n}_{t}$. The unitary transformation $\hat{V}_{\mathrm{invert}}$ does not change the Josephson energy term either. Since the negative sign can be absorbed into the definition of the cavity annihilation and creation operators, the spectrum of the transmon-cavity system is invariant under the sign inversion of the offset charge. From the periodicity and the symmetry, it is enough to investigate the region $0\leq N_{g}\leq 0.5$ for the identification.
@@
Figure 5: Photon-number dependence of the cavity frequency with the offset charges from 00 to $0.5$ for (a) the ground-state ladder |$\ket{\overline{g},\overline{n}}$ and (b) the excited-state ladder |$\ket{\overline{e},\overline{n}}$. We use $(E_{C}/\hbar\omega_{c},E_{J}/\hbar\omega_{c},g/\omega_{c})=(5.0\times 10^{-2},1.6,2.5\times 10^{-2})$ for the system parameters. For the cases $N_{g}=0.1$ in the ground-state ladder and $N_{g}=0.3$ in the excited-state ladder, the energy window $\delta$ is set to $1.5\times 10^{-2}\hbar\omega_{c}$. For the other cases, we set $\delta$ to $1.0\times 10^{-2}\hbar\omega_{c}$.
@@
Figure 5 shows the photon-number dependence of the cavity frequency with the offset charges from 0 to 0.5 for the ground- (excited-)state ladder |$\ket{\overline{g},\overline{n}}$ (|$\ket{\overline{e},\overline{n}}$). For the ground-state ladder |$\ket{\overline{g},\overline{n}}$, the cavity frequencies for the different offset charges do not show visible difference up to $n\sim 150$ apart from the resonant peaks. The little dependence on the offset charge implies that the weights of higher transmon excited states in the labeled states |$\ket{\overline{g},\overline{n}}$ are small for $n\lesssim 150$ (apart from the resonant point). The location of the resonant point strongly depends on the offset charge, as expected. The resonant point can fluctuate between $n\sim 15$ and $n\sim 130$ without controlling the offset charge. Therefore, a transition to a higher excited state may take place with a relatively small photon number $n\sim O(10)$.
@@
For the excited-state ladder |$\ket{\overline{e},\overline{n}}$, the cavity frequencies show visible difference for $n\gtrsim 20$. This implies that the weights of higher transmon excited states are not negligible in this region. For the case with $N_{g}=0.5$, the resonant peak is located around $n\sim 10$. Consequently, a transition to a higher excited state may take place with a relatively small photon number also in the case of the labeled states |$\ket{\overline{e},\overline{n}}$. In this way, the photon-number dependence of the cavity frequency obtained by the labeling method provides the information on photon number where the transitions leading to leakage from the computational space can occur.
VI Summary
@@
In this paper, we proposed a new method to label the eigenstates of a qubit-cavity system based on the continuity of qubit occupancy. Starting from the eigenstate without photon occupation, the system is expected to follow the labeled states under adiabatic injection of photons into the cavity. Besides, the effective resonant frequency of the cavity with photon occupation can be estimated from the eigenenergies of the labeled states. Our proposed method finds candidates for the next labeled state based on their eigenenergies. Then, one selects the next labeled state that makes the change of the qubit occupancy as small as possible. Comparing the obtained photon-number dependence of the cavity frequency and the evolution under cavity driving, we confirmed that our proposed scheme can be applied to a broader parameter range compared to the previous approach [21].